OpenStudy (nuccioreggie):

The amounts of household waste generated in a region during two years were 23,400,000 tons and 48,100,000 tons. What was the total household waste generated over these two years in the region, expressed in scientific notation? 7.15 ⋅ 107 tons 7.15 ⋅ 1014 tons 71.5 ⋅ 107 tons 71.5 ⋅ 1014 tons

OpenStudy (nuccioreggie):

@mathmath333

OpenStudy (nuccioreggie):

@Nnesha

OpenStudy (nuccioreggie):

@satellite73

OpenStudy (nuccioreggie):

Please help dont leave

OpenStudy (nuccioreggie):

@pooja195

OpenStudy (nuccioreggie):

@whpalmer4

OpenStudy (whpalmer4):

Add the two numbers together, what do you get?

OpenStudy (nuccioreggie):

I got 71,500,000

OpenStudy (whpalmer4):

Good. How do you write that in scientific notation?

OpenStudy (nuccioreggie):

I got B

OpenStudy (whpalmer4):

\[7.15*10^{14}\]? You sure about that? \[10^{14} = 100,000,000,000,000\]

OpenStudy (nuccioreggie):

Wait no D i jus got them mixed up its D

OpenStudy (whpalmer4):

Scientific notation has two parts, mantissa and exponent. The mantissa is that number out from, and it is always between 1 and 10 (or -1 and -10). Come on, stop guessing. D is\[71.5*10^{14} = 71.5*100,000,000,000,000\]

OpenStudy (whpalmer4):

\[10^0=1\]\[10^1=10\]\[10^2=100\]\[10^3=1,000\]\[10^4=10,000\]\[10^5=100,000\]\[10^6=1,000,000\]\[10^7=10,000,000\]etc.

OpenStudy (nuccioreggie):

Ok sorry i just did math on paper and got A

OpenStudy (whpalmer4):

that's more like it!

OpenStudy (nuccioreggie):

Thanks can u help me with a few more

OpenStudy (whpalmer4):

sure, fire away, got a couple minutes